### v2

@@ 1.4 Other Related Work
Fair division of indivisible goods is a vast area of research in recent years, a survey paper by Amanatidis et al., “Fair Division of Indivisible Goods: Recent Progress and Open Questions”, [7], includes 12 pages of references, and the authors stress that this is not an exhaustive survey and refer to yet other surveys (e.g., [36] with a more economic perspective) that focus on specific aspects of the problem.
@@
Fair division notions can be roughly partitioned into share based [9] and envy based [17]. In share based fair division one ignores what the others get, one is happy if one gets value above some threshold (that does not depend on others, only on their cardinality), proportionality and MMS are share based. In envy based fair division one decides if one is happy or not depending on the others’ allocated bundles. The basic envy based fairness notion is that being envy free (one does not seek to exchange allocations with the other). For indivisible goods, envy freeness is impossible in general and thus relaxations were proposed. These include envy freeness up to one good (EF1) [35, 21] and envy freeness up to any good (EFX) [23]. EF1 allocations are known to exist, the question of when do EFX allocations exist is a notoriously hard open problem. The number of citations dealing with envy free, EF1 and EFX allocations is quite daunting, see [7]. In this paper we deal with share based, and more specifically MMS fairness.
Besides MMS and $\alpha$-MMS, other relaxations/approximations to proportionality include proportionality up to one good (PROP1), Round Robin Share, Pessimistic proportional share [24], proportional up to any good (PROPx) [8] which need not exist, and proportionality up to the maximin item (PROPm) [13] which are known to exist for 5 agents. Allocations that are both PROP1 and Pareto efficient exist [24].
@@
Prior to Budish [21], Hill also considered the maximin share [32] and gave lower bounds on the maximin share as a function of the item values.
There was a long series of papers improving upon results for $\alpha$-MMS, complexity improvements and improvements in the quality of the result [37, 5, 34, 29, 14], $\alpha=3/4$ was achieved in [30, 28, 3] and finally [1] improved upon $3/4$ by a small constant.
A formal definition of shares and of feasible shares appears in [9, 11], MMS itself is not feasible as it may not exist, even for additive valuations. A recent share-based notion is quantile-share [12], which is guaranteed to exist for additive valuations and, more generally, for monotone valuations (then called universally feasible), assuming the Erdös Matching Conjecture holds.
The work of [2] studies EF1 and Pareto optimal allocations of chores under chore duplication, whereas our work focuses on chore deletion. In a similar spirit to our resource augmentation approach, [41] and [15] investigate the computational complexity of achieving envy-free and proportional allocations when a limited number of goods can be divided among multiple agents.
@@
There is a long history of considering resource augmentation in various forms in computer science and economics, see chapter on resource augmentation in “Beyond the Worst-Case Analysis of Algorithms” [39]. Resource augmentation appears in Bulow Klemperer [22] where simple auctions with one extra bidder guarantee revenue at least as large as that of complex auctions without the extra bidder. This was further studied in [31] and since then has been used in many contexts such as auction design, for both revenue and welfare, prophet inequalities, etc. It also goes by the name of competition complexity, see [26, 40, 10, 19, 20, 25].
@@ 2 Model and Preliminaries
Following standard usage, define $[k]:=\{1,2,\ldots,k\}$ for integer $k\geq 1$.
An instance of a fair division problem is given by a tuple $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, where $M$ is a set of $m$ indivisible goods, and $N$ is a set of $n$ agents. Every agent $i$ has a valuation function $v_{i}:2^{M}\to\mathbb{R}_{\geq 0}$, associating a non-negative value to every set $S\subseteq M$ of goods, denoted by $v_{i}(S)$. As standard, we assume that $v_{i}$ is monotone, i.e., for any $T\subseteq S\subseteq M$, $v_{i}(T)\leq v_{i}(S)$, and normalized, so that $v_{i}(\emptyset)=0$.
A valuation function $v_{i}$ is said to be additive if for every set of items $S\subseteq M$, $v_{i}(S)=\sum_{g\in S}v_{i}(\{g\})$. We abuse the notation and write $v_{i}(g)$ or $v_{ig}$ instead of $v_{i}(\{g\})$ for $g\in M$.
An allocation of the set $M$ is a tuple $(A_{1},\dots,A_{n})$, where $A_{i}$ denotes the bundle allocated to agent $i$, such that $\bigcup_{i\in[n]}A_{i}=M$ (i.e., all items are allocated), and $A_{i}\cap A_{j}=\emptyset$ for any $i\neq j$ (i.e., no item is allocated to more than one agent). We denote the collection of all allocations of the set $M$ among $n$ agents by $\mathcal{A}_{n}(M)$.
The maximin share (MMS) of agent $i$ [21] is the value that agent $i$ can guarantee by splitting the set $M$ into $n$ bundles, getting the worst out of them. Formally:
Definition 1 (Maximin share (MMS)).
For an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ with $n$ agents, the maximin share (MMS) of agent $i$ with valuation function $v_{i}$, denoted by $\mu_{i}^{n}(M)$, is given by
$\mu_{i}^{n}(M)=\max_{A\in\mathcal{A}_{n}(M)}\min_{S\in A}v_{i}(S).$
@@
Without loss of generality, we remove all agents whose MMS value is 00. These agents get their full MMS with empty bundles. As valuations can be scaled, we may assume, without loss of generality that the MMS value for every agent is one. For additive valuations, this implies that every agent values the grand bundle (all items) at $n$ or more.
Any partition that realizes the MMS value of agent $i$, is said to be an MMS partition of agent $i$. That is, a partition $P=(P_{1},\ldots,P_{n})$ of the goods into $n$ bundles is an MMS partition of agent $i$, if and only if $\mu_{i}^{n}(M)=\min_{j}v_{i}(P_{j})$.
@@
Definition 2 ($\alpha$-MMS).
Given a fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and $\alpha>0$, an allocation $A$ is $\alpha$-MMS if and only if for all agents $i\in N$, $v_{i}(A_{i})\geq\alpha\cdot\mu_{i}^{n}(M)$.
@@
Valid reductions allow us to remove a subset of agents and items while ensuring that the MMS values of the remaining agents do not decrease. The following lemma is due to [6, 16].
Lemma 1 ([6, 16]).
@@
By applying Lemma 1 repeatedly, we conclude that for any $k<\min\{m,n\}$, the MMS of agent $i$ may only increase if we remove an arbitrary set of $k$ goods and an arbitrary set of $k$ agents. This is shown in the following corollary.
Corollary 1.
@@
An allocation with copies is an allocation where goods may be duplicated, with different instances of a same good allocated to different agents. An agent may receive at most one instance of each good. An allocation with copies is a relaxation of an allocation where the two or more bundles may have a non-empty intersection.
Definition 3 ($(t,k)$-allocation).
Given a fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, a $(t,k)$-allocation $A=(A_{1},\dots,A_{n})$ is a collection of sets of items, such that
@@
    The total number of extra copies is at most $t$, i.e., $\sum_{i\in N}|A_{i}|\leq|M|+t$.
@@
    The number of extra copies of any good $g\in M$ is at most $k$, i.e., for every $g\in M$, |$|\{i\in N\mid g\in A_{i}\}|\leq k+1$.
Note that, the definition above implies that no agent is allowed to receive more than a single instance of each good.
We also consider valid reductions with copies which appear in Section 5.2.
3 MMS with Copies: General Valuations
In this section we provide (tight) bounds on the number of copies needed for full MMS under general valuations. Section 3.1 establishes bounds on the total number of copies, while Section 3.2 gives bounds on the number of copies per good.
3.1 A Tight Bound of $m/e$ on the Total Number of Copies
In this section we give a tight bound on the total number of copies required to guarantee original MMS value to all agents under monotone valuations.
Our lower bound is based on a combinatorial construction which considers $m=n^{n}$ goods positioned on the $n$-dimensional cube, so each good may be identified with $n$ coordinates, $(d_{1},\dots,d_{n})$. The valuations are such that each of the agent is almost “single-minded”, with agent $i$ only interested in bundles that contain all $n^{n-1}$ goods that share the same value in the $d_{i}$ coordinate. Those $n$ different bundles (one for each possible value of $d_{i}$), make up her MMS partition. A simple counting argument shows that the only way to satisfy all agents simultaneously is to copy $(n-1)^{n}$ goods, which is a $1/e$-fraction of the total number of goods, as $n$ tends to infinity.
We complement this result by showing, using the probabilistic method, that in every instance there exists an allocation which satisfies all agents, and copies at most $\left(1-\frac{1}{n}\right)^{n}$-fraction of the goods.
Theorem 1.
For any $n$, there exists an instance with $n$ monotone valuations such that the total number of copies needed is $\left(\frac{n-1}{n}\right)^{n}$-fraction of the goods, and there is at least one good which must be copied $n-1$ times.
@@ 3.1 A Tight Bound of $m/e$ on the Total Number of Copies
Each agent $i\in[n]$ is interested in one of $n$ disjoint bundles, $P_{i}^{1},\dots,P_{i}^{n}$ of $n^{n-1}$ goods each, where $P_{i}^{j}=\{g_{d_{1},\dots,d_{n}}\mid d_{i}=j\}$. Formally, agent $i$’s valuation is monotone and for every $S\subseteq M$ it is defined to be:
$v_{i}(S)=\begin{cases}1&\exists j\in[n]\text{ s.t. }P_{i}^{j}\subseteq S\\ 0&\text{otherwise}\end{cases}$
Agent $i$’s MMS partition is $(P_{i}^{1},\dots,P_{i}^{n})$ and her MMS value is 1.
Observe that for any $k$ agents, $i_{1},\dots,i_{k}\in[n]$ and for any $k$ indexes $j_{1},\dots,j_{k}\in[n]$, it holds that
$|P_{i_{1}}^{j_{1}}\cap P_{i_{2}}^{j_{2}}\cap\dots\cap P_{i_{k}}^{j_{k}}|=|\{g_{d_{1},\dots,d_{n}}\mid\forall l\in[k]\;\;d_{i_{l}}=j_{l}\}|=n^{n-k}.$
If we wish to find an allocation with copies $A=(A_{1},\dots,A_{n})$ such that for every $i$, it holds that $v_{i}(A_{i})\geq 1$, while minimizing the number of items we copy, then it is without loss to assume that $A_{i}=P_{i}^{j_{i}}$ for some $j_{i}\in[n]$. Thus, the total number of goods (copies included) is $\sum_{i\in[n]}|A_{i}|=\sum_{i\in[n]}|P_{i}^{j_{i}}|=n\cdot n^{n-1}=n^{n}$. The number of different goods used is, using the inclusion-exclusion principle:
$\left|\bigcup_{i\in[n]}P_{i}^{j_{i}}\right|$ $=$ $\sum_{k=1}^{n}(-1)^{k+1}\binom{n}{k}n^{n-k}=n^{n}-n^{n}-\left(\sum_{k=1}^{n}(-1)^{k}\binom{n}{k}n^{n-k}\right)$
@@
which implies that the total number of copies is $(n-1)^{n}$. Additionally, in any such allocation there exists a good which is allocated $n$ times, and thus duplicated $n-1$ times, as $g_{j_{1},j_{2},\dots,j_{n}}\in\bigcap_{l=1}^{n}P_{i_{l}}^{j_{l}}$, and the claim follows. ∎
Below we prove a matching upper bound on the number of total copies required. We show that there always exists an allocation which assigns to every agent $i$ one of his MMS bundles and makes no more than $\left(1-\frac{1}{n}\right)^{n}\cdot m$ additional copies.
Proposition 1.

### v3

@@ 1.3 Other Related Work
Fair division of indivisible goods is a vast area of research in recent years, a survey paper by Amanatidis et al., “Fair Division of Indivisible Goods: Recent Progress and Open Questions”, [Amanatidis_2023], includes 12 pages of references, and the authors stress that this is not an exhaustive survey and refer to yet other surveys (e.g., [moulinsurvey] with a more economic perspective) that focus on specific aspects of the problem.
@@
Fair division notions can be roughly partitioned into share based [BabaioffF22] and envy based [Brams_Taylor_95]. In share based fair division one ignores what the others get, one is happy if one gets value above some threshold (that does not depend on others, only on their cardinality), proportionality and MMS are share based. In envy based fair division one decides if one is happy or not depending on the others’ allocated bundles. The basic envy based fairness notion is that being envy free (one does not seek to exchange allocations with the other). For indivisible goods, envy freeness is impossible in general and thus relaxations were proposed. These include envy freeness up to one good (EF1) [Lipton2004, budish2011combinatorial] and envy freeness up to any good (EFX) [Caragiannis19]. EF1 allocations are known to exist, the question of when do EFX allocations exist is a notoriously hard open problem. The number of citations dealing with envy free, EF1 and EFX allocations is quite daunting, see [Amanatidis_2023]. In this paper we deal with share based, and more specifically MMS fairness.
@@
Prior to Budish [budish2011combinatorial], Hill also considered the maximin share [Hill87] and gave lower bounds on the maximin share as a function of the item values.
There was a long series of papers improving upon results for $\alpha$-MMS for additive valuations, complexity improvements and improvements in the quality of the result [Amanatidis_2017, Kurokawa18, garg2019approximating, barman2020approximation], which led to $\alpha=3/4$ [ghodsi2022fair, garg2021improved, akrami2023simplification], and slightly beyond [akrami2023breaking34barrierapproximate]. Recently, [heidari2025improved1013] proved the existence of a $\frac{10}{13}$-MMS, which was then improved to $\frac{7}{9}$-MMS by [huang2025fptas79].
$\alpha$-MMS allocations have also been studied in the settings with more general valuations such as submodular [barman2020approximation, ghodsi2022fair, uziahu2023, chekuri20241, chekuri2025cycle, uziahu2023], XOS [ghodsi2022fair, SEDDIGHIN2024104049, Akrami2023], and subadditive [SEDDIGHIN2024104049, feige2025, seddighin2025beating, feige2025fair]. Notably, and as discussed in [seddighin2025beating], one can deduce from Lemma 6.1 in [SEDDIGHIN2024104049] the existence of an allocation in which every item is copied at most $\log_{3/2}n+1$ times and guarantees $1/4$-MMS for subadditive valuations. Interestingly, this “multi-allocation”, as it is termed in [SEDDIGHIN2024104049, seddighin2025beating], is then used to construct an approximate MMS allocation without copies, a process was later improved by [feige2025multi]. This exemplifies how the resource augmentation approach is not only of independent interest, but can be used as a stepping stone to improve results in the non-augmented problem.
Besides MMS and $\alpha$-MMS, other relaxations/approximations to proportionality include proportionality up to one good (PROP1), Round Robin Share, Pessimistic proportional share [Conitzerprop1], proportional up to any good (PROPx) [aziz2020polynomial] which need not exist, and proportionality up to the maximin item (PROPm) [baklanov2021achievingproportionalitymaximinitem] which are known to exist for 5 agents. Allocations that are both PROP1 and Pareto efficient exist [Conitzerprop1]. A formal definition of shares and of feasible shares appears in [BabaioffF22, BabaioffEF24], MMS itself is not feasible as it may not exist, even for additive valuations. A recent share-based notion is quantile-share [BabichenkoFHN24], which is guaranteed to exist for additive valuations and, more generally, for monotone valuations (then called universally feasible), assuming the Erdös Matching Conjecture holds. Recently, Feige [feige2025residualmaximinshare] introduced a new feasible relaxation of MMS called Residual MMS. Finally, the work of [gafni2023unified] studies envy-freeness (and relaxations) for a setting in which multiple copies of each good exist, and each agent can get at most one copy of a good.
@@
There is a long history of considering resource augmentation in various forms in computer science and economics, see chapter on resource augmentation in “Beyond the Worst-Case Analysis of Algorithms” [Roughgarden_2021]. Resource augmentation appears in Bulow Klemperer [bulowklemperer96] where simple auctions with one extra bidder guarantee revenue at least as large as that of complex auctions without the extra bidder. This was further studied in [HartlineR09] and since then has been used in many contexts such as auction design, for both revenue and welfare, prophet inequalities, etc. It also goes by the name of competition complexity, see [EdenFFTW17a, RoughgardenTY20, babaioff2019welfaremaximization, BrustleCDV24, Brustle0DEFV24, derakhshan2024settlingcompetitioncomplexity].
@@ 2 Model and Preliminaries
In this section we present our model and preliminaries for MMS allocations of goods.
Following standard usage, define $k]:=\{1,2,\ldots,k\}$ for integer $k\geq 1$. An instance of a fair division problem is given by a tuple $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$, where $M$ is a set of $m$ indivisible goods, and $N$ is a set of $n$ agents. Every agent $i$ has a valuation function $v_{i}:2^{M}\to\mathbb{R}_{\geq 0}$, associating a non-negative value to every set $S\subseteq M$ of goods, denoted by $v_{i}(S)$. For subsets $S\subseteq M$ and $g\in M$ we will write $S+g$ to denote the union $S\cup\{g\}$. As standard, we assume that $v_{i}$ is monotone, i.e., for any $T\subseteq S\subseteq M$, $v_{i}(T)\leq v_{i}(S)$, and normalized, so that $v_{i}(\emptyset)=0$.
In this work we also consider instances where the valuation functions are of a specific class. A valuation function $v_{i}$ is said to be additive if for every set of items $S\subseteq M$, $v_{i}(S)=\sum_{g\in S}v_{i}(\{g\})$. In this case, we abuse the notation and write $v_{i}(g)$ or $v_{ig}$ instead of $v_{i}(\{g\})$ for $g\in M$.
An allocation of the set $M$ is a tuple $A=(A_{1},\dots A_{n})$, where $A_{i}$ denotes the bundle allocated to agent $i$, such that $\bigcup_{i\in[n]}A_{i}=M$ (i.e., all items are allocated), and $A_{i}\cap A_{j}=\emptyset$ for any $i\neq j$ (i.e., no item is allocated to more than one agent). We denote the collection of all allocations of the set $M$ among $n$ agents by $\Pi_{n}(M)$.
The maximin share (MMS) [budish2011combinatorial] of agent $i$ is the value that agent $i$ can guarantee by splitting the set $M$ into $n$ bundles, getting the worst out of them. Formally,
Definition 2.1 (Maximin Share (MMS)).
For an instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ with $n$ agents, the maximin share (MMS) of agent $i$ with valuation function $v_{i}$, denoted by $\mu_{i}^{n}(M)$, is given by
$\mu_{i}^{n}({M})=\max_{(A_{1},\ldots,A_{n})\in\Pi_{n}(M)}\ \min_{1\leq\ell\leq n}v_{i}\left(A_{\ell}\right).$
@@
Without loss of generality, we remove all agents whose MMS value is 0. These agents get their full MMS with empty bundles. As valuations can be scaled, we may assume, without loss of generality that the MMS value for every agent is one. For additive valuations, this implies that every agent values the grand bundle (all items) at $n$ or more.
Any partition that realizes the MMS value of agent $i$, is said to be an MMS partition of agent $i$. That is, a partition $P=(P_{1},\dots,P_{n})$ of the goods into $n$ bundles is an MMS partition of agent $i$, if and only if $\mu_{i}^{n}({M})=\min_{j}v_{i}(P_{j})$.
@@
Definition 2.2 ($\alpha$-MMS).
Given a fair division instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and $\alpha>0$, an allocation $A$ is $\alpha$-MMS if and only if for all agents $i\in N$, $v_{i}(A_{i})\geq\alpha\cdot\mu_{i}^{n}({M})$.
@@
Valid reductions allow us to remove a subset of agents and items while ensuring that the MMS values of the remaining agents do not decrease. The following lemma is due to [amanatidis2018comparing, bouveret2016characterizing].
Lemma 2.3 ([amanatidis2018comparing, bouveret2016characterizing]).
@@
By applying Lemma 2.3 repeatedly, we conclude that for any $k<\min\{m,n\}$, the MMS of agent $i$ may only increase if we remove an arbitrary set of $k$ goods and an arbitrary set of $k$ agents. This is shown in the following corollary.
Corollary 2.4.
@@
An allocation with copies is an allocation where goods may be duplicated, with different instances of the same good allocated to different agents. An agent may receive at most one instance of each good. An allocation with copies is a relaxation of an allocation where the two or more bundles may have a non-empty intersection.
Definition 2.5 ($(t,k)$-allocation).
In a fair division instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$, a collection of subsets of items $A=(A_{1},\dots A_{n})$ is said to be a $(t,k)$-allocation if
@@
    For each good $g\in M$, the number of assigned extra copies is at most $k$. That is, for each $g\in M$ and (copy count) $k_{g}\coloneqq\max\Big\{0,\ \ \left|\left\{i\in N\mid g\in A_{i}\right\}\right|-1\Big\}$, it holds that $k_{g}\leq k$.
@@
    The total number of extra copies is at most $t$, i.e., $\sum_{g\in M}k_{g}\leq t$.
Note that, the definition above implies that no agent is allowed to receive more than a single instance of each good. We also consider valid reductions with copies which appear in Section 4.2. Also, for a $(t,k)$-allocation $A=(A_{1},\dots A_{n})$ it holds that $\sum_{i\in N}|A_{i}|\leq|M|+t$.
Classes of Valuation Functions.
We consider different classes of valuation functions $v:2^{M}\to\mathbb{R}_{+}$, defined below.
  • Additive: For any $S\subseteq M$, it holds that $v(S)=\sum_{g\in S}v(\{g\})$.
  • Submodular: For any $S,T\subseteq M$, $v(S\cup T)\leq v(S)+v(T)-v(S\cap T)$.
  • XOS: There exists a finite set of additive functions $\mathcal{L}$ such that for any $S\subseteq M$, $v(S)=\max_{\ell\in\mathcal{L}}\ell(S)$.
  • Subadditive: For any $S,T\subseteq M$, $v(S\cup T)\leq v(S)+v(T)$.
It is well known that $\text{Additive}\subsetneq\text{Submodular}\subsetneq\text{XOS}\subsetneq\text{Subadditive}$ [lehmann2001combinatorial].
3 Exact Maximin Shares
3.1 XOS Valuations and Beyond
In this section we show that for any instance with general monotone valuation there exists an MMS allocation using at most $m/e$ total copies and no more than $\ln m/\ln\ln m$ copies of an individual good. We show that these bounds are tight even when all valuations are XOS (a strict subclass of subadditive valuations). In Section 3.1.1 we describe our lower bound construction, in which $m/e$ total copies and $\ln m/\ln\ln m$ copies of a single good are required for achieving an MMS allocation. In sections 3.1.2 and 3.1.3 we prove the matching upper bounds for the total and individual number of copies, respectively.
[…]
Our lower bound is based on a combinatorial construction which considers $m=n^{n}$ goods positioned on the $n$-dimensional cube, so each good may be identified with $n$ coordinates, $(d_{1},\dots,d_{n})$. The valuations are such that the agent $i$’s unique MMS partition is such that MMS bundle $j$, $P_{i}^{j}$, contains all $n^{n-1}$ goods that satisfy $d_{i}=j$. The value of a certain set of goods $S$ is determined by the maximal intersection $|S\cap P_{i}^{j}|$. A simple counting argument shows that the only way to satisfy all agents simultaneously is to copy $(n-1)^{n}$ goods, which is a $1/e$-fraction of the total number of goods, as $n$ tends to infinity.
Theorem 3.1.
For any $n$, there exists an instance with $n$ XOS valuations such that, the total number of copies required to achieve MMS, is a $\left(\frac{n-1}{n}\right)^{n}$-fraction of the goods, and there is at least one good which must be copied (n1)(n-1) times.
@@ 3.1 XOS Valuations and Beyond
We design agent $i$’s valuation so that her MMS partition, $P^{i}=(P^{i}_{1},\dots,P^{i}_{n})$, is such that $P^{i}_{j}=\{g_{d_{1},\dots,d_{n}}\mid d_{i}=j\}$ and $v_{i}(P^{i}_{j})=1$ for any $j\in[n]$. Observe that each MMS bundle has exactly $n^{n-1}$ goods.
The valuation of agent $i$ is defined using $n$ additive functions, $f_{i}^{j}$ for any $j\in[n]$, where a good has positive value if only if its $i$th coordinate is $j$ (i.e., it belongs to $j$th MMS bundle of agent $i$), namely $f_{i}^{j}(g_{d_{1},\dots,d_{n}})=\frac{1}{n^{n-1}}\cdot I[d_{i}=j]$. We define $v_{i}(S)=\max_{j\in[n]}f_{i}^{j}(S)$, which makes $v_{i}$ an XOS function. Another way to write $v_{i}$ is $v_{i}(S)=\frac{1}{n^{n-1}}\cdot\max_{j\in[n]}|S\cap P^{i}_{j}|$, which immediately implies the following observation.
Observation 3.2.
For any agent $i\in[n]$, $v_{i}(S)=1$ if and only if there exists $j\in[n]$ such that $P^{i}_{j}\subseteq S$.
Observe that for any $k$ distinct agents, $i_{1},\dots,i_{k}\in[n]$ and for any $k$ indices $j_{1},\dots,j_{k}\in[n]$,
$|P^{i_{1}}_{j_{1}}\cap P^{i_{2}}_{j_{2}}\cap\dots\cap P^{i_{k}}_{j_{k}}|=|\{g_{d_{1},\dots,d_{n}}\mid\forall l\in[k]\;\;d_{i_{l}}=j_{l}\}|=n^{n-k}.$
If we wish to find an allocation with copies $A=(A_{1},\dots A_{n})$ such that for every $i$, it holds that $v_{i}(A_{i})\geq 1$, while minimizing the number of items we copy, then by 3.2 it is without loss of generality to assume that $A_{i}=P^{i}_{j_{i}}$ for some $j_{i}\in[n]$. Thus, the total number of goods (copies included) is $\sum_{i\in[n]}|A_{i}|=\sum_{i\in[n]}|P^{i}_{j_{i}}|=n\cdot n^{n-1}=n^{n}$. The number of different goods used is, using the inclusion-exclusion principle:
$\left|\bigcup_{i\in[n]}P^{i}_{j_{i}}\right|$ $=$ $\sum_{k=1}^{n}(-1)^{k+1}\binom{n}{k}n^{n-k}=n^{n}-n^{n}-\left(\sum_{k=1}^{n}(-1)^{k}\binom{n}{k}n^{n-k}\right)$
@@
which implies that the total number of copies is $(n-1)^{n}$. Additionally, in any such allocation there exists a good which is allocated $n$ times, and thus duplicated n1n-1 times, as $g_{j_{1},j_{2},\dots,j_{n}}\in\bigcap_{l=1}^{n}P_{j_{l}}^{i_{l}}$, and the claim follows. ∎
Remark.
The above construction implies that under XOS valuations, one cannot hope to achieve a 1-out-of-$f(n)$ MMS, for any function $f(n)$ of the number of agents, see 4.
3.1.2 Upper Bound on the Total Number of Copies for General Monotone Valuations
Below we prove a matching upper bound on the number of total copies required. We show that there always exists an allocation which assigns to every agent one of her MMS bundles and makes no more than $\left(1-\frac{1}{n}\right)^{n}\cdot m$ additional copies.
Proposition 3.3.